# Computational bottlenecks for denoising diffusions

### v1

@@ A.1 Properties of the estimator $\hat{\boldsymbol{m}}(\,\cdot\,)$
The proof of this proposition is a modification of the one in [CLR17], and will be presented in Appendix C. Note that Proposition A.1 directly implies the first inequality in Condition M2v of Theorem 3, as $\|\hat{\boldsymbol{m}}(\boldsymbol{y}_{t},t)-\boldsymbol{x}\|\leq 2$.
@@
In the proof of Theorem 3, we will also make use of the following estimates, whose proof is deferred to the Appendix D.
Lemma A.3.
@@
A.2 Analysis of the diffusion process: Proof of Theorem 3
@@
Note that Propositions A.1, A.2 hold under these choices, and in particular $\boldsymbol{g}_{t}\sim{\mathsf{N}}({\boldsymbol{0}},t{\boldsymbol{I}}_{n\times n})$ at all times. Also the sequence of random vectors $\boldsymbol{g}_{\ell\Delta}$, $\ell\in{\mathbb{N}}$ can be generated via $\boldsymbol{g}_{\ell\Delta}=\boldsymbol{g}_{(\ell-1)\Delta}+\sqrt{\Delta}\hat{\boldsymbol{g}}_{\ell}$, for some i.i.d. standard normal vectors $\{\hat{\boldsymbol{g}}_{\ell}\}_{\ell\geq 0}$.
@@
Before proving this proposition, let us show that it implies Condition M3v of Theorem 3. Indeed we claim that, with high probability, for all $\ell\in{\mathbb{N}}$, $\hat{\boldsymbol{m}}(\hat{\boldsymbol{y}}_{\ell\Delta},\ell\Delta)={\boldsymbol{0}}$ and $\hat{\boldsymbol{y}}_{\ell\Delta}=\boldsymbol{z}_{\ell{\Delta}}$. This is proven by induction over $\ell$. Indeed, if it holds up to a certain $\ell-1\in{\mathbb{N}}$, then we have $\hat{\boldsymbol{y}}_{\ell\Delta}=\boldsymbol{z}_{\ell\Delta}$ by Eq. (41) whence it follows that $\boldsymbol{A}_{t,+}=\boldsymbol{w}_{t}/\sqrt{t}$, for $t=\ell\Delta$ (c.f. Algorithm 2, line 4) and therefore $\hat{\boldsymbol{m}}(\hat{\boldsymbol{y}}_{t},t)={\boldsymbol{0}}$ by Proposition A.4 (because the condition in Algorithm 2, line 6, is never passed).
@@
This concludes the proof of Theorem 3. We next turn to proving Proposition A.4.
Proof of Proposition A.4.
@@
We have from [BVH16], for every $x\geq 0$:
@@
From Lemma A.3, we obtain that
@@
Remark A.1.
We remark that Assumption (C2) in the proof above is technically not needed, meaning that the additional noise stream $\boldsymbol{g}_{t}$ can in fact be discarded entirely: an appropriate thresholding of $\boldsymbol{v}_{t}$, the top eigenvector of $\eta_{s}(\boldsymbol{A}_{t,+})$, as in Algorithm 1, will also suffice to satisfy all conditions of Theorem 3, although $\boldsymbol{x}$ will not be recovered exactly; some $o(k)$ positions outside the support of $\boldsymbol{x}$ will also be chosen, at most. The reason for this is that the alignment $|\langle\boldsymbol{v}_{t},\boldsymbol{u}\rangle|=1-o_{n}(1)$ already, from a close inspection of our proof of Proposition A.1. Regarding the proof of Proposition A.4 above, one can easily realize that even if $\varepsilon=0$, it will go through without any modification. We choose to keep our formulation of Algorithm 2 as faithful to the original work of [CLR17] as possible to discuss a variety of approaches, and leave this to the interested reader.
Appendix B Proof of Proposition 2.1
@@
It is clear that as $\ell\to\infty$, this quantity converges to 00.
Appendix C Proof of Proposition A.1
@@
for some constant $C>0$ that we choose such that $\max_{ij\leq n}|Z_{ij}|\leq C\sqrt{\log n}$ with probability at least $1-n^{-D}$ ($D$ is defined in the statement of Proposition A.1). We have $\tilde{\boldsymbol{Z}}=\eta_{s}(\boldsymbol{Z})$. By [BVH16], there exists an absolute constant $c>0$ such that for every $u>0$:
@@
Appendix D Proof of Lemma A.3
Proof.
[…]
Appendix E Auxiliary lemmas for Section 5
E.1 Proof of Lemma 5.1
@@ Appendix E Auxiliary lemmas for Section 5
E.2 Proof of Lemma 5.3
Using Lemma 5.2, we can take $x=n^{-1/4}$, say, and $\theta=\sqrt{1+\delta}$ for $\delta\geq n^{-c_{0}}$ for some small enough $c_{0}>0$, to get that
@@
Appendix F Proof of Proposition 6.1
@@
We conduct our analysis conditional on $\boldsymbol{u}$. Let $\boldsymbol{v}_{t}$ be a top eigenvector of $\boldsymbol{A}_{t}$. For $t=(1+\delta)n/2$, $|\langle\boldsymbol{v}_{t},\boldsymbol{u}\rangle|\overset{a.s.}{\to}\sqrt{\delta/(1+\delta)}$, so that with high probability $|\langle\boldsymbol{v}_{t},\boldsymbol{u}\rangle|\geq\sqrt{\delta}/(2\sqrt{1+\delta})$. If $t\geq(1+\delta)n/2$, we can use Fact F.2 to obtain the same result. By choosing $\varepsilon$ such that $2\varepsilon<\sqrt{\delta/(1+\delta)}$, we know from standard concentration of the alignment (Lemma 6.6) that $|\langle\boldsymbol{v}_{t},\boldsymbol{u}\rangle|\geq 2\varepsilon$ with probability at least $1-\exp(-cn)$ for some $c>0$ possibly dependent on $(\varepsilon,\delta)$.
@@
Appendix G Proof of Lemma 6.3
@@
Appendix H Proof of Lemma 6.4
@@
As with the proof of Proposition 6.1, we first deal with the denominator $\|\boldsymbol{g}\|$: indeed, sub-exponential concentration gives us
@@
Appendix I Proof of Lemma 6.5
@@
$\sup_{t_{\ell}\leq t\leq t_{\ell+1}}\left|\boldsymbol{v}_{t}^{{\sf T}}\boldsymbol{W}_{t_{i}}\boldsymbol{v}_{t}-\lambda_{1}(\boldsymbol{W}_{t_{\ell}})\right|\leq 2\|\boldsymbol{W}_{t}-\boldsymbol{W}_{t_{\ell}}\|_{\text{op}}\leq 16\sqrt{(t_{\ell+1}-t_{\ell})n}$
@@
Appendix J Proof of Lemma 6.6
By Weyl’s inequality, $\boldsymbol{W}\mapsto\lambda_{1}(\boldsymbol{Y})$ (with $\boldsymbol{Y}=\theta\boldsymbol{v}\boldsymbol{v}^{{\sf T}}+\boldsymbol{W}$) is a 1-Lipschitz function and therefore, by Borell inequality (and [BBAP05]), letting $\lambda_{*}(\theta):=\theta+1/\theta$, for any $\varepsilon>0$,
@@
Appendix K Proofs of reduction results
K.1 Proof of Theorem 5
@@
(See for instance [MW23] or [AMS23] for examples of this calculation.) Let now $\boldsymbol{y}_{T}\stackrel{{\scriptstyle{\rm d}}}{{=}}T\boldsymbol{x}+\sqrt{T}\boldsymbol{g}^{\prime}$ for $(\boldsymbol{x},\boldsymbol{g}^{\prime})\sim\mu\otimes{\mathsf{N}}(\boldsymbol{0},\boldsymbol{I})$. Another application of Assumption $(a)$ implies that this can be coupled to $\hat{\boldsymbol{y}}^{*}_{T}$ so that $\mathbb{E}\|\boldsymbol{y}_{T}-\hat{\boldsymbol{y}}^{*}_{T}\|\leq\varepsilon$, and therefore

### v2

@@ D.1 Properties of the estimator $\hat{\boldsymbol{m}}(\,\cdot\,)$
The proof of this proposition is a modification of the one in Cai et al. [2017], and will be presented in Appendix F. Note that Proposition D.1 directly implies the first inequality in Condition M2v of Theorem 3, as $\|\hat{\boldsymbol{m}}(\boldsymbol{y}_{t},t)-\boldsymbol{x}\|\leq 2$.
@@
In the proof of Theorem 3, we will also make use of the following estimates, whose proof is deferred to the Appendix I.
Lemma D.3.
@@
D.2 Analysis of the diffusion process: Proof of Theorem 3
@@
Note that Propositions D.1, D.2 hold under these choices, and in particular $\boldsymbol{g}_{t}\sim{\mathsf{N}}({\boldsymbol{0}},t{\boldsymbol{I}}_{n\times n})$ at all times. Also the sequence of random vectors $\boldsymbol{g}_{\ell\Delta}$, $\ell\in{\mathbb{N}}$ can be generated via $\boldsymbol{g}_{\ell\Delta}=\boldsymbol{g}_{(\ell-1)\Delta}+\sqrt{\Delta}\hat{\boldsymbol{g}}_{\ell}$, for some i.i.d. standard normal vectors $\{\hat{\boldsymbol{g}}_{\ell}\}_{\ell\geq 0}$.
@@
Before proving this proposition, let us show that it implies Condition M3v of Theorem 3. Indeed we claim that, with high probability, for all $\ell\in{\mathbb{N}}$, $\hat{\boldsymbol{m}}(\hat{\boldsymbol{y}}_{\ell\Delta},\ell\Delta)={\boldsymbol{0}}$ and $\hat{\boldsymbol{y}}_{\ell\Delta}=\boldsymbol{z}_{\ell{\Delta}}$. This is proven by induction over $\ell$. Indeed, if it holds up to a certain $\ell-1\in{\mathbb{N}}$, then we have $\hat{\boldsymbol{y}}_{\ell\Delta}=\boldsymbol{z}_{\ell\Delta}$ by Eq. (42) whence it follows that $\boldsymbol{A}_{t,+}=\boldsymbol{w}_{t}/\sqrt{t}$, for $t=\ell\Delta$ (c.f. Algorithm 2, line 4) and therefore $\hat{\boldsymbol{m}}(\hat{\boldsymbol{y}}_{t},t)={\boldsymbol{0}}$ by Proposition D.4 (because the condition in Algorithm 2, line 6, is never passed).
@@
This concludes the proof of Theorem 3. We next turn to proving Proposition D.4.
Proof of Proposition D.4.
@@
We have from Bandeira and van Handel [2016], for every $x\geq 0$:
@@
From Lemma D.3, we obtain that
@@
We remark that Assumption (C2) in the proof above is technically not needed, meaning that the additional noise stream $\boldsymbol{g}_{t}$ can in fact be discarded entirely: an appropriate thresholding of $\boldsymbol{v}_{t}$, the top eigenvector of $\eta_{s}(\boldsymbol{A}_{t,+})$, as in Algorithm 1, will also suffice to satisfy all conditions of Theorem 3, although $\boldsymbol{x}$ will not be recovered exactly; some $o(k)$ positions outside the support of $\boldsymbol{x}$ will also be chosen, at most. The reason for this is that the alignment $|\langle\boldsymbol{v}_{t},\boldsymbol{u}\rangle|=1-o_{n}(1)$ already, from a close inspection of our proof of Proposition D.1. Regarding the proof of Proposition D.4 above, one can easily realize that even if $\varepsilon=0$, it will go through without any modification. We choose to keep our formulation of Algorithm 2 as faithful to the original work of Cai et al. [2017] as possible to discuss a variety of approaches, and leave this to the interested reader.
Appendix E Proof of Proposition 2.1
@@
It is clear that as $\ell\to\infty$, this quantity converges to 00. Hence we are done.
Appendix F Proof of Proposition D.1
@@
for some constant $C>0$ that we choose such that $\max_{ij\leq n}|Z_{ij}|\leq C\sqrt{\log n}$ with probability at least $1-n^{-D}$ ($D$ is defined in the statement of Proposition D.1). We have $\tilde{\boldsymbol{Z}}=\eta_{s}(\boldsymbol{Z})$. By Bandeira and van Handel [2016], there exists an absolute constant $c>0$ such that for every $u>0$:
@@
Appendix G Auxiliary lemmas for Section C
G.1 Proof of Lemma C.1
@@ Appendix G Auxiliary lemmas for Section C
G.2 Proof of Lemma C.3
Using Lemma C.2, we can take $x=n^{-1/4}$, say, and $\theta=\sqrt{1+\delta}$ for $\delta\geq n^{-c_{0}}$ for some small enough $c_{0}>0$, to get that
@@
Appendix H Proof of Proposition B.1
@@
We conduct our analysis conditional on $\boldsymbol{u}$. Let $\boldsymbol{v}_{t}$ be a top eigenvector of $\boldsymbol{A}_{t}$. For $t=(1+\delta)n/2$, $|\langle\boldsymbol{v}_{t},\boldsymbol{u}\rangle|\overset{a.s.}{\to}\sqrt{\delta/(1+\delta)}$, so that with high probability $|\langle\boldsymbol{v}_{t},\boldsymbol{u}\rangle|\geq\sqrt{\delta}/(2\sqrt{1+\delta})$. If $t\geq(1+\delta)n/2$, we can use Fact H.2 to obtain the same result. By choosing $\varepsilon$ such that $2\varepsilon<\sqrt{\delta/(1+\delta)}$, we know from standard concentration of the alignment (Lemma B.6) that $|\langle\boldsymbol{v}_{t},\boldsymbol{u}\rangle|\geq 2\varepsilon$ with probability at least $1-\exp(-cn)$ for some $c>0$ possibly dependent on $(\varepsilon,\delta)$.
@@
Appendix I Proof of Lemma D.3
Proof.
[…]
Appendix J Proof of Lemma B.3
@@
Appendix K Proof of Lemma B.4
@@
As with the proof of Proposition B.1, we first deal with the denominator $\|\boldsymbol{g}\|$: indeed, sub-exponential concentration gives us
@@
Appendix L Proof of Lemma B.5
@@
$\sup_{t_{\ell}\leq t\leq t_{\ell+1}}\left|\boldsymbol{v}_{t}^{{\sf T}}\boldsymbol{W}_{t_{\ell}}\boldsymbol{v}_{t}-\lambda_{1}(\boldsymbol{W}_{t_{\ell}})\right|\leq 2\sup_{t_{\ell}\leq t\leq t_{\ell+1}}\|\boldsymbol{W}_{t}-\boldsymbol{W}_{t_{\ell}}\|_{\text{op}}\leq 32\sqrt{(t_{\ell+1}-t_{\ell})n}$
@@
Appendix M Proof of Lemma B.6
By Weyl’s inequality, $\boldsymbol{W}\mapsto\lambda_{1}(\boldsymbol{Y})$ (with $\boldsymbol{Y}=\theta\boldsymbol{v}\boldsymbol{v}^{{\sf T}}+\boldsymbol{W}$) is a 1-Lipschitz function and therefore, by Borell inequality (and Baik et al. [2005]), letting $\lambda_{*}(\theta):=\theta+1/\theta$, for any $\varepsilon>0$,
@@
Appendix N Proofs of reduction results
N.1 Proof of Theorem 4
N.2 Proof of Theorem 5
@@
(See for instance Montanari and Wu [2023] or Alaoui et al. [2023] for examples of this calculation.) Let now $\boldsymbol{y}_{T}\stackrel{{\scriptstyle{\rm d}}}{{=}}T\boldsymbol{x}+\sqrt{T}\boldsymbol{g}^{\prime}$ for $(\boldsymbol{x},\boldsymbol{g}^{\prime})\sim\mu\otimes{\mathsf{N}}(\boldsymbol{0},\boldsymbol{I})$. Another application of Assumption $(a)$ implies that this can be coupled to $\hat{\boldsymbol{y}}^{*}_{T}$ so that $\mathbb{E}\|\boldsymbol{y}_{T}-\hat{\boldsymbol{y}}^{*}_{T}\|\leq\varepsilon$, and therefore